OpenStudy (anonymous):

someone please explain in simple steps how to do this?Solve the system. y = -3x + 5 and y = 2x + 5 (0, 5) (5, 0) (-5/2, 0) no solution infinitely many solutions

OpenStudy (compassionate):

You would plug the points in given to the equations. If the statement is true with both of them, then choose that point. If they are all correct, than infinite solutions. If none of them are correct, then no solutions.

OpenStudy (anonymous):

is it 0,5?

OpenStudy (mathstudent55):

One method of solving a system of equations is called substitution. You solve one equation for one variable, and then you substitute what that variable is equal to everywhere you see it in the other equation. Since in this system of equations, both equations are already solved for one variable, y, it makes this an ideal situation to use substitution. y = -3x + 5 y = 2x + 5 Take what y equals in the first equation, -3x + 5, and insert it for y in the second equation. -3x + 5 = 2x + 5 Now solve for x: -5x = 0 x = 0 Once you have x, substitute what x is in one of the original equations and solve for y: y = 2x + 5 y = 2(0) = 5 y = 0 + 5 y = 5 Solution: (0, 5) Your answer is correct.

OpenStudy (anonymous):

thank you so much i know that took a while to write!

OpenStudy (mathstudent55):

You're welcome. The important thing is if it helped you.

OpenStudy (anonymous):

it did thanks!

OpenStudy (mathstudent55):

Great. You're welcome.